A janitor invested $11,000 for one year, part at 8%, and the rest at 11%. The 11% investment returns $450 more than the 8% investment. How much was invested at 11%?
Answer by stanbon(26285)

A janitor invested $11,000 for one year, part at 8%, and the rest at 11%. The 11% investment returns $450 more than the 8% investment. How much was invested at 11%?
---------------
Let amount invested at 8% be "x".
Then amount invested at 11% is "11,000-x"
----------------------------------------------
Equation:
0.11(11000-x)-0.08x = 450
11(11000-x) - 8x = 45000
121000 - 11x - 8x = 45000
-19x = -76000
x = $4000 (amt invested at 8%
11000-x = $7000 (amt invested at 11%)
=========================================
Cheers,
Stan H.